OpenStudy (anonymous):

Roma drove 270 mi from her hometown to the city. During her return trip, she was able to increase her speed by 9 mph. If her return trip took 1 hr less time, find her original speed.

OpenStudy (anonymous):

that's 2430

OpenStudy (anonymous):

pleeease hellp!! :)

OpenStudy (anonymous):

never mind use a equation d=rt

OpenStudy (anonymous):

distance= rate times time

OpenStudy (anonymous):

so the distance is 270 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i found out that the orig. and return distances are 270, the orig. rate is x, the return rate is x+9, the orig. time is 120/x and the return time is 120/x+6 i just don't know what to do w/ the numbers

OpenStudy (anonymous):

pleasee heelllpp!!! i only have 1 more try on my homework haha

OpenStudy (anonymous):

Let s = speed to the city then (s+9) = speed home : Write a time equation, time = dist/speed : To city time - return time = 1 hr less

OpenStudy (anonymous):

Multiply by s(s+9), results: 270(s+9) - 270s = 1s(s+9) :

OpenStudy (anonymous):

270s +2430 - 270s= 1s(s+9)

OpenStudy (anonymous):

i got s^2+9s-2430=0

OpenStudy (anonymous):

you must distribute 270 to 9 and s

OpenStudy (anonymous):

i did but then I did like terms

OpenStudy (anonymous):

here use this link it show the exact question but with different numbers it should guide you: http://www.algebra.com/algebra/homework/word/travel/Travel_Word_Problems.faq.question.472436.html

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

tell me what you get ok?

OpenStudy (anonymous):

i got 45 mph :)

OpenStudy (anonymous):

im soo sorry i have to go im heading to a wedding!

OpenStudy (anonymous):

its okay thanks so much!